A stock's returns have the following distribution:

Demand for the
Company's Products
Probability of This
Demand Occurring
Rate of Return If
This Demand Occurs
Weak 0.1 (22%)
Below average 0.1 (13)   
Average 0.3 17  
Above average 0.4 32  
Strong 0.1 73  
1.0

Assume the risk-free rate is 4%. Calculate the stock's expected return, standard deviation, coefficient of variation, and Sharpe ratio. Do not round intermediate calculations. Round your answers to two decimal places.

Stock's expected return: ____%

Standard deviation: ____%

Coefficient of variation: ____

Sharpe ratio: ____
Answer #1

Expected return =0.1*-22%+0.1*-13%+0.3*17%+0.4*32%+0.1*73% =21.70%
Standard Deviation =(0.1*(-22%-21.70%)^2+0.1*(-13%-21.70%)^2+0.3*(17%-21.70%)^2+0.4*(32%-21.70%)^2+0.1*(73%-21.70%)^2)^0.5 =24.9722% or 24.97%

Coefficient of variation =Standard Deviation/Expected Return =24.9722%/21.70% =1.15
Sharpe Ratio =(Expected return -Risk Free rate)/Standard Deviation =(21.70%-4%)/24.9722%=0.71